OpenStudy (anonymous):

How to simplify (x^2y^3)^4(xy^4)^-3/(x^2y)

OpenStudy (anonymous):

some one please help me...i have my exam in 2hours..

OpenStudy (anonymous):

\[\frac{\left(x^2 y^3\right)^4}{\left(x y^4\right)^3 \left(x^2 y\right)}=\frac{x^8 y^{12}}{x^5 y^{13}}=\frac{x^3}{y} \]